A university found that 10% of its students withdraw without completing the introductory statistics course. Assume that 12 students registered for the course. What is the probability that at least 2 students will withdraw from the course? (use the binomial distribution table)

Group of answer choices

0.2389

0.2301

0.3890

0.3410
Answer #1

X ~ Bin ( n , p)

Where n = 12 , p = 0.10

Binomial probability distribution is

P(X) = nCx * px * ( 1 - p)n-x

P(X >= 2) = 1 - P(X <= 1)

= 1 - [ P(X = 0) + P(X = 1) ]

= 1 - [ 12C0 * 0.100 * (1 - 0.10)12 + 12C1 * 0.101 * (1 - 0.10)11 ]

= 0.3410